# Volatility Parametrizations with Random Coefficients: Analytic Flexibility for Implied Volatility Surfaces

### v2

@@ 3.1 Randomized Flat Volatility
We obtain the discretized randomization price $V_{c/p}(T,K;(\mu,\nu),N_{q})$ for $N_{q}$ quadrature points by substituting the integral in the equation with a finite sum [^6] Note that this a similar result as Brigo and Mercurio  [4], since the randomized price is a mixture of Black-Scholes prices.
@@
The quadrature points $\sigma_{n}$ and quadrature weights $\lambda_{n}$ can be obtained by computing the moments $\mu_{i}={\mathbb{E}[\vartheta^{i}]={\rm e}^{i\mu+i^{2}\nu^{2}/2}}$ for all $i\leq 2N$ and computing the matrices $M,R$ and $J$ as described in A. To obtain the implied volatility surface of the randomized parametrization, we can use a root-finding algorithm to find the implied volatility for each option price given by $V_{c/p}(T,K;(\mu,\nu),N_{q})$, or we can derive the expansion terms from Equation 2.19 to obtain an analytical expression for the implied volatility surface $\hat{\sigma}(T,K;(\mu,\nu))$. We examine a slice of the randomized surface for two different sets of parameters. Figure 3 show the results for the expansions with different numbers of coefficients and the “exact” implied volatility obtained through a root-solving algorithm. In the experiment, we use $r=2\%,T=2$ and $N_{q}=4$ quadrature points.
@@
The randomization of the previously flat volatility surface yielded a new parametrization with two parameters $\mu,\nu$, which is able to form the characteristic “volatility smile”. The result is a variant of the lognormal mixture model [4], which is known to be able to form a volatility smile. The main difference between the approaches is that the randomization defines the volatility smile using only two parameters, $\mu$ and $\eta$ (excluding the choice of $N_{q}=4$), offering a more parsimonious alternative to the lognormal mixture model, which uses seven parameters. While this may limit some flexibility compared to the discrete mixture model, it simplifies calibration and reduces computational complexity.
Figure 3 shows that the 6th-order approximation is able to closely match the exact implied volatility obtained from the root-finding algorithm. We observe that the fit is worse on the right-hand side for the lower-order expansions, although the shape of the implied volatility smile is similar. The reason is that the overall volatility level is lower on the right-hand side. This apparent increase in the importance of the higher-order coefficients can be explained by examining Equation 2.19, where each term is divided by a power of $a_{0}=\frac{1}{2}\sigma(0)\sqrt{T}$. It should be reiterated, however, that there is no theoretical limit to the amount of expansion terms to be calculated.
The advantage of using the $n$-order expansion versus the application of a root-finding algorithm to the randomized pricing surface is based on the computation complexity. We examine the computational time for an increasing number of pairs $(T,K)$. The root finding algorithm uses Brent’s method and terminates once the relative difference between two estimates is at most $081e-$ 08. If available, the method uses the implied volatility from the previous (usually neighboring) pair of $(T,K)$ to increase the efficiency. Table 1 shows the comparison of the 2nd, 4th and 6th order approximation to the Brent’s method. The analytical method demonstrates a negligible increase in computation time as the number of strike/time pairs increases, due to its computational efficiency and independence from iterative procedures required by root-finding algorithms like Brent’s method.
Number of strike/time $10^{3}$ $10^{4}$ $5.0\cdot 10^{4}$ $10^{5}$
@@
We conclude that the randomization of the flat volatility surface using a lognormal random variable allows us to create volatility smiles by adding a single parameter. Although the shape of the smile can be altered by changing the parameter or even using an alternative distribution for $\vartheta$, the additional flexibility is limited (see also [4] for a discussion on this issue).
@@ 3.2 Randomized SABR parametrization
In the second example, we consider the SABR parametrization, a well-known volatility parametrization introduced by Hagan et al. [15]. The parametrization is derived from the SABR model [^7] The SABR model is a stochastic volatility model with the same parameters, such that ${\rm d}F_{t}$ $=\sigma_{t}F_{t}^{\beta}{\rm d}W_{t},$ ${\rm d}\sigma_{t}$ $=\gamma\sigma_{t}{\rm d}Z_{t},$ with $\sigma_{0}=\alpha$ and the stochastic drivers such that ${\rm d}\langle W_{t},Z_{t}\rangle=\rho{\rm d}t$. and given by the formula
@@
Although the parametrization defines an entire volatility surface, market practice is to use the SABR parametrization “slice-wise”, meaning that the calibration is done per volatility slice $\{K\mapsto\hat{\sigma}_{T_{n}}(K;\overline{p}),n\leq N\}$ for the set of $N$ expiries observed in the market. The surface is then constructed by a linear interpolation, which is free of arbitrage under certain conditions. The details on the interpolation and the conditions are provided later.
@@
where $\Gamma(k)$ is the Gamma function. The random variable is almost surely positive, which makes it suitable for a randomization of $\gamma$, since the domain of $\gamma$ is $[0,\infty)$. In this case, the randomized price function is given by
@@
Figure 4: Implied volatility skew: The parameters $k$ and $\theta$ have a strong effect on the shape of the implied volatility.
@@
By randomizing $\gamma$, we parameters that modulate not only the level of the skew, but also the curvature of the shape around the ATM point. To evaluate the impact of this modification, we conduct the following experiment: using the randomized SABR parametrization, we generate two sets of implied volatility quotes across 40 strikes. We then apply an optimizer to fit the traditional SABR parametrization to these quotes, resulting in nearly identical parameter values for both sets, specifically $(\beta,\alpha,\rho,\gamma)=(0.9,0.25,-0.135,3.5)$.
@@
The experiment in Figure 5 shows that the SABR parametrization cannot account for the detailed curvature. The parametrization is limited by the choice of a single parameter. The randomization of the parameter helps to gain flexiblity, which can be essential in fitting the market.
@@ 3.3 Fitting the model to the market
To study the flexibility of the new randomized SABR parametrization, we apply it to real data to examine its fit. We collect a set of short-maturity index options on the SPX index and obtain the market quotes for European put and call options on the 31st of July 2024 for a range of strikes and expiries. The data is downloaded from the Cboe data shop [5], which is based on actual transactions occurring on the Cboe exchange. Table 2 summarizes the data of the option chains and expiries available in the set. When both put and call quotes are available we select the most liquid quote per strike based on the open interest, which are the out-the-money quotes.
Expiry Date Days to Maturity N Quotes Min Strike Max Strike
Aug 16 2024 16 381 3175 6900
Sep 20 2024 51 356 3100 7200
Oct 18 2024 79 340 3100 9000
Nov 15 2024 107 289 3025 8600
Table 2: Data summary implied volatilities quotes on Jul 31st, 2024
For the randomized SABR parametrization, we choose $N_{q}=2$ quadrature points. Additionally, we fix $\beta=0.9$ to reduce the number of parameters to calibrate, following common practice and empirical observations that this value provides a good balance between model flexibility and calibration stability [20]. The remaining parameters are fitted to the market quotes using an optimization scheme. The optimization can be achieved on the randomized prices given by Equation 3.9 or on the implied volatility surface using the expansion of Theorem 2.4. Table 3 shows the calibrated parameters for the randomized SABR parametrization.
Expiry ${\beta}$ $\alpha$ $\rho$ $k$ $\theta$ Var $\Gamma$
Aug 0.9 0.335 -0.7 1.775 1.378 3.371
Sep 0.9 0.319 -0.681 3.872 0.455 0.802
Oct 0.9 0.318 -0.674 3.032 0.446 0.603
Nov 0.9 0.338 -0.687 4.916 0.271 0.361
@@
To compare the fit, we also run the optimization on the regular SABR parametrization as a reference. The fit of the two models is shown in Figure 6 compared to the market quotes. We see that the randomized SABR parametrization has an excellent fit and can replicate the skew observed on the short-term maturity options on the SPX index.
Sum Squared Errors (SSE) Mean Squared Errors (MSE)
Month Regular SABR Randomized SABR Regular SABR Randomized SABR
Aug 0.1248596 0.0414336 3.53E-04 1.17E-04
Sep 0.0192344 0.0009572 5.92E-05 2.90E-06
Oct 0.0233494 0.0113573 7.56E-05 3.68E-05
Nov 0.015732 0.0006516 5.74E-05 2.40E-06
Figure 6: Fit to the market: The randomized SABR parametrization has an excellent fit to the market data.
We observe from Table 3 that the variance of the randomized parameter $\vartheta$, which is given by $\mathbb{V}\text{ar}(\vartheta)=k\theta^{2}$ decreases as the time to maturity increases. This shows that the longer the time-to-maturity, the less randomization is required to fit this observed market regime.
Since we fit the SABR parametrization slice-wise, it remains to show the construction of an arbitrage-free surface with the slices. As the slices are necessarily arbitrage-free, the construction must be free of calendar arbitrage, governed by Condition v) from 2.1. The condition states that the pricing function must be non-decreasing in the time direction, meaning that an interpolation scheme in the time direction is sufficient for the construction. Suppose that we calibrate a set of slices $\{\hat{\sigma}_{T_{i}}(K;\overline{p}_{i}^{*},N_{q}):i\leq N\}$, which are arbitrage-free in strike direction, and suppose that there is no calendar-spread arbitrage in the market. This means that, if $T_{i}\leq T_{j}$, then
@@
It is well-known [9] that the absence of calendar-spread arbitrage is equivalent to the condition that the total implied variance is increasing, i.e.
@@
Under this condition, a linear interpolation in total implied variance suffices to obtain an arbitrage-free volatility surface from the slices. Let $a\in[0,1]$ be such that $T=(1-a)T_{i}+aT_{j}$ and define
@@
This shows that if the calibrated slices are consistent in time, meaning that Equation 3.10 is true, the linear interpolation and the calibrated slices define an arbitrage-free surface for all $T\in[T_{1},T_{N}]$. This concludes the randomization of the SABR parametrization.
@@ 4 Randomized Spot Volatility Parametrizations & Near Expiry Options for Equities
In this section of this paper, we consider an extension of the randomization of parametric volatility surfaces to a randomization of the spot price $S_{0}$, which follows the same principles as the regular randomization. The new randomization, which we refer to as the spot randomization, replaces the usually predetermined variable of the spot price $S_{0}$ with a random variable $\vartheta$. We will show below that, under certain conditions, such a randomization of the spot price is arbitrage-free, and that this parametrization is particularly effective to model option markets of very short-term options, known as near-expiry options. Near expiry options, also known as zero-day-expiry options (0DTE) options [^8] To be precise, 0DTE options have a time-to-maturity of less than a day. Since we also consider options of 1-3 days expiry, we prefer the expression near expiry options. are options whose maturity date is imminent, i.e., only a few days in the future, or even expire at the end of the day. On the day of an earning announcement, these options chains sometimes exhibit volatility slices with one or more concave sections, such as the W-shaped implied volatility slice or the “mustache”- shape [1, 10]. The unusual shape of implied volatility is an indication of a bimodal or multimodal risk-neutral probability distribution [^9] Although note that this is neither a sufficient nor a necessary condition., i.e. a probability distribution function which exhibits two modes. The rationale is that the uncertainty from the earnings yields a multi-modal risk-neutral probability distribution of $S_{T}$, reflecting the evolution of the stock price given different scenarios from the earnings (i.e., positive surprise vs negative surprise). Traditional implied volatility parametrizations, which have their origins in stochastic diffusion models, such as the SABR parametrization or the SVI, struggle to produce such shapes since the stochastic driver, which is a diffusion process, is inherently single-modal. The novel randomization method on the spot price offers a valuable solution in this case, as the randomization of the spot price offers an effective way to create superpositions of the PDFs with different (single) modes, yielding a mixture density of a similar kind to the lognormal mixture model as described in Glasserman and Pirjol [10]. We will prove that the randomized volatility shape is arbitrage-free and show that it produces excellent results on the empirical data, utilizing data from the ticker AMZN on the day of the earning announcement for Q1 in April 2018. Since the options we consider are short-expiry options, we will focus the fitting of the data on a single volatility slice for a fixed time $T$. Although the randomization still provides an entire implied volatility surface, the main purpose of the spot randomization is to fit the non-standard volatility shape and to provide a viable price for any strike for the given expiry.
4.1 Randomized Spot Prices

### v3

@@ 3.1 Randomized Flat Volatility
We obtain the discretized randomization price $V_{c/p}(T,K;(\mu,\nu),N_{q})$ for $N_{q}$ quadrature points by substituting the integral in the equation with a finite sum [^7] Note that this a similar result as Brigo and Mercurio  [brigo2002lognormal], since the randomized price is a mixture of Black-Scholes prices.
@@
The quadrature points $\sigma_{n}$ and quadrature weights $\lambda_{n}$ can be obtained by computing the moments $\mu_{i}={\mathbb{E}[\vartheta^{i}]={\rm e}^{i\mu+i^{2}\nu^{2}/2}}$ for all $i\leq 2N$ and computing the matrices $M,R$ and $J$ as described in A. To obtain the implied volatility surface of the randomized parametrization, we can use a root-finding algorithm to find the implied volatility for each option price given by $V_{c/p}(T,K;(\mu,\nu),N_{q})$, or we can derive the expansion terms from Equation 2.21 to obtain an analytical expression for the implied volatility surface $\hat{\sigma}(T,K;(\mu,\nu))$. We examine a slice of the randomized surface for two different sets of parameters. Figure 3 show the results for the expansions with different numbers of coefficients and the “exact” implied volatility obtained through a root-solving algorithm. In the experiment, we use $r=2\%,T=3$ and $N_{q}=4$ quadrature points.
@@
The randomization of the previously flat volatility surface yielded a new parametrization with two parameters $\mu,\nu$, which is able to form the characteristic “volatility smile”. The result is a variant of the lognormal mixture model [brigo2002lognormal], which is known to be able to form a volatility smile. The main difference between the approaches is that the randomization defines the volatility smile using only two parameters, $\mu$ and $\eta$ (excluding the choice of $N_{q}=4$), offering a more parsimonious alternative to the lognormal mixture model, which uses 7 parameters [^8] Four volatilities and four weights, where one of the weights is given by the condition that they sum to 1.. While this may limit some flexibility compared to the discrete mixture model, it simplifies calibration and reduces computational complexity.
Figure 3 shows that the 6th-order approximation is able to closely match the exact implied volatility obtained from the root-finding algorithm. We observe that the fit is worse on the right-hand side for the lower-order expansions, although the shape of the implied volatility smile is similar. The reason is that the overall volatility level is lower on the right-hand side. This apparent increase in the importance of the higher-order coefficients can be explained by examining Equation 2.21, where each term is divided by a power of $\Sigma_{0}=\frac{1}{2}P_{(T,K)}(0)\sqrt{T}$.
Remark (Radius of Convergence).
The Taylor expansion of implied volatility around $m=0$ is subject to a possible finite radius of convergence, beyond which the series may diverge even for an infinite number of expansion terms. While the exact radius is parameter-dependent and theoretically inaccessible, our numerical experiments in Figure 3 indicate that the approximation converges within the plotted range. For deeper OTM options, the accuracy of the expansion should be compared to the implied volatilities obtained from a root solving algorithm.
The advantage of using the $n$-order expansion versus the application of a root-finding algorithm to the randomized pricing surface is based on the computation complexity. We examine the computational time for an increasing number of pairs $(T,K)$. The root finding algorithm uses Brent’s method and terminates once the relative difference between two estimates is at most 1e081e-08. If available, the method uses the implied volatility from the previous (usually neighboring) pair of $(T,K)$ to increase the efficiency. Table 1 shows the comparison of the 2nd, 4th and 6th order approximation to the Brent’s method. The analytical method demonstrates a negligible increase in computation time as the number of strike/time pairs increases due to its computational efficiency and independence from iterative procedures required by root-finding algorithms like Brent’s method.
Number of $(T,K)$ $10^{3}$ $10^{4}$ $5.0\cdot 10^{4}$ $10^{5}$
@@
We conclude that the randomization of the flat volatility surface using a lognormal random variable allows us to create volatility smiles by adding a single parameter. Although the shape of the smile can be altered by changing the parameter or even using an alternative distribution for $\vartheta$, the additional flexibility is limited (see also [brigo2002lognormal] for a discussion on this issue).
@@ 3.2 Randomized SABR parametrization
In the second example, we consider the SABR parametrization, a well-known volatility parametrization introduced by Hagan et al. [Hagan:2002]. The parametrization is derived from the SABR model [^9] The SABR model is a stochastic volatility model with the same parameters, such that ${\rm d}F_{t}$ $=\sigma_{t}F_{t}^{\beta}{\rm d}W_{t},$ ${\rm d}\sigma_{t}$ $=\gamma\sigma_{t}{\rm d}Z_{t},$ with $\sigma_{0}=\alpha$ and the stochastic drivers such that ${\rm d}\langle W_{t},Z_{t}\rangle=\rho{\rm d}t$. and given by the formula
@@
Although the parametrization defines an entire volatility surface, market practice is to use the SABR parametrization “slice-wise”, meaning that the calibration is done per volatility slice $\{K\mapsto\hat{\sigma}_{T_{n}}(K;\overline{p}),n\leq N\}$ for the set of $N$ expiries observed in the market. The surface is then constructed by a linear interpolation, which is not guaranteed to be free of arbitrage. The details on the interpolation and the conditions are provided later.
@@
where $\Gamma(k)$ is the Gamma function. The random variable is almost surely positive, which makes it suitable for a randomization of $\gamma$, since the domain of $\gamma$ is $0,\infty)$. In this case, the randomized price function is given by
@@
Figure 4: Implied volatility skew: The parameters $k$ and $\theta$ have a strong effect on the shape of the implied volatility. Left: $\left(\beta,\rho,\theta\right)=(0.9,-0.7,1)$, Right: $\left(\beta,\rho,k\right)=(0.9,-0.7,1)$. The parameter $\alpha$ is chosen such that the ATM-vol equals $0.2$. For both plots, we chose $N_{q}=10$.
@@
By randomizing $\gamma$, we add parameters that modulate not only the level of the skew but also the curvature of the shape around the ATM point. To evaluate the impact of this modification, we conduct the following experiment: using the randomized SABR parametrization, we generate two sets of implied volatility quotes across 40 strikes with parameters $(\beta,\alpha,\rho,k,\theta)=(0.9,0.03,-0.1,0.5,2)$ and (0.9,0.03,0.1,0.1,10)(0.9,0.03,-0.1,0.1,10) and $N_{q}=10$. We then apply an optimizer to fit the traditional SABR parametrization to these quotes, resulting in nearly identical optimal parameter values for both sets, specifically $(\beta,\alpha,\rho,\gamma)=(0.9,0.25,-0.135,3.5)$.
@@
The experiment in Figure 5 shows that the SABR parametrization cannot account for the detailed curvature produced by the randomized SABR. The parametrization is limited by the choice of a single parameter. The randomization of the parameter helps to gain flexibility, which can be essential in fitting the market, as we will see in the next subsection.
@@ 3.3 Fitting the model to the market
To study the flexibility of the new randomized SABR parametrization, we apply it to real data to examine its fit. We collect a set of option quotes of index options on the SPX index and obtain the market prices for European put and call options on the 31st of July 2024 for a range of strikes and expiries. The data is downloaded from the Cboe data shop [cboeSPX], which is based on actual transactions occurring on the Cboe exchange. Information about the data and how it is processed is found in 6.
For the randomized SABR parametrization, we choose $N_{q}=2$ quadrature points and choose a Gamma distribution for $\vartheta$ as described before. We fix $\beta=0.9$ to reduce the number of parameters to calibrate, following common practice and empirical observations that this value provides a good balance between model flexibility and calibration stability [Oosterlee_Grzelak_2020]. The remaining parameters are fitted to the market quotes using an optimization scheme. The optimization can be achieved on the randomized prices given by Equation 3.9 or on the implied volatility surface using the expansion of Theorem 2.5. Table 2 shows the calibrated parameters for the randomized SABR parametrization. Since we choose $N_{q}=2$, the surface is a mixture of two price surfaces with different $\theta_{n}$ and we report the respective values to show the impact of the Gamma parameter $k,\theta$ on the mixture parameters.
Remark (Choice of $N_{q}$).
Generally, a larger choice of $N_{q}$ means that the discretized randomization approximates the continuous randomization better, and $N_{q}=2$ does not seem like a good approximation. However, since the discretization is arbitrage-free for any $N_{q}\in\mathbb{N}$, an exact approximation of a continuous $\vartheta$ is not necessary, and $N_{q}$ can be treated simply as a parameter to further specify the nature of the randomization of the parameters. In this case, $N_{q}=2$ offers an excellent fit.
Expiry Randomization Parameters Var Mixture Parameters
$\beta$} $\alpha$ $\rho$ $k$ $\theta$ Var $\vartheta$ $\theta_{1}$ $\theta_{2}$ $\lambda_{1}$ $\lambda_{2}$
Aug 24 0.9 0.322 -0.595 2.379 1.04 2.572 1.602 5.424 0.772 0.228
Sep 24 0.9 0.308 -0.617 10.587 0.187 0.37 1.531 2.804 0.647 0.353
Oct 24 0.9 0.306 -0.63 11.724 0.145 0.247 1.33 2.366 0.64 0.36
Nov 24 0.9 0.325 -0.649 9.987 0.144 0.206 1.102 2.054 0.651 0.349
Dec 24 0.9 0.329 -0.657 14.387 0.09 0.116 1.031 1.737 0.627 0.373
Jan 25 0.9 0.328 -0.673 15.439 0.078 0.094 0.968 1.601 0.623 0.377
Feb 25 0.9 0.332 -0.69 15.401 0.071 0.079 0.883 1.461 0.623 0.377
Mar 25 0.9 0.335 -0.689 20.794 0.051 0.053 0.865 1.337 0.607 0.393
Jun 25 0.9 0.341 -0.695 16.435 0.053 0.046 0.702 1.144 0.62 0.38
Sep 25 0.9 0.347 -0.681 15.715 0.048 0.037 0.609 1.004 0.622 0.378
Dec 25 0.9 0.35 -0.656 15.016 0.045 0.031 0.543 0.904 0.625 0.375
Dec 26 0.9 0.359 -0.599 460.433 0.001 0.001 0.508 0.558 0.523 0.477
Dec 27 0.9 0.377 -0.595 681.713 0.001 0 0.398 0.43 0.519 0.481
@@
To compare the fit, we run the optimization on the regular SABR parametrization and a lognormal mixture (according to [brigo2002lognormal]) as a benchmark. For the lognormal mixture we use 4 mixture terms and a displacement parameter to obtain the best possible fit. This yields an 8-parameter optimization problem (displacement, 4 volatilities, 3 weights), which we fit using a global optimizer. The fit of the two models is shown in Figure 6 compared to the market quotes for the short time-to-maturities and in Figure 7 for the long time-to-maturities. Table 3 shows the observed mean-squared errors of the implied volatilities over the entire dataset. We see that the randomized SABR parametrization has an excellent fit and can replicate the skew observed in the short-term maturity options on the SPX index. To show the consistency of the improved fit over time we repeat the experiment for the option chains at the two consecutive months ends - 2024-08-30 and 2024-09-30. The additional fitting results are presented in 6.
Figure 6: Short time to maturity fit: the graph shows the fit of the randomized SABR vs the benchmarks for short time-to-maturities.
Figure 7: Long time to maturity fit: the graph shows the fit of the randomized SABR vs the benchmarks for long time-to-maturities.
Jul 31 2024 Mean Squared Errors (MSE)
Expiry Month Lognormal Mixture Regular SABR Randomized SABR
Aug 24 7.24E-04 4.82E-04 1.23E-05
Sep 24 8.26E-04 1.00E-04 6.23E-05
Oct 24 6.73E-04 6.74E-05 4.69E-05
Nov 24 6.33E-04 6.88E-05 2.80E-05
Dec 24 6.58E-04 6.69E-05 1.23E-05
Jan 25 7.61E-04 8.02E-05 1.01E-05
Feb 25 1.08E-03 8.90E-05 9.10E-06
Mar 25 1.54E-03 8.76E-05 7.10E-06
Jun 25 1.65E-03 1.31E-04 3.20E-06
Sep 25 9.14E-04 1.61E-04 1.00E-06
Dec 25 7.00E-04 1.73E-04 1.10E-06
Dec 26 7.07E-04 3.32E-04 1.20E-05
Dec 27 3.68E-04 2.74E-04 3.29E-05
Table 3: Mean Square Error across entire dataset
We observe from Table 2 that the variance of the randomized parameter $\vartheta$, which is given by $\mathbb{V}\text{ar}(\vartheta)=k\theta^{2}$ decreases as the time to maturity increases. This shows that the longer the time-to-maturity, the less randomization is required to fit this observed market regime. While $\theta_{1}$ and $\theta_{2}$ are far apart for the short-expiry months, the values are much closer for longer time-to-maturities. However, from the graphs it is also noticeable that the regular SABR in unable to fit the price surface well.
Since we fit the SABR parametrization slice-wise, the construction of an entire surface requires combining the slices into a unified surface. This is a common practice for the SABR parametrization, and is commonly solved by using a linear interpolation in the total implied variance [bender2020arbitrage]. The interpolation guarantees the absence of calendar arbitrage, but not absence of butterfly arbitrage within the interpolated slices. To illustrate this, suppose that we calibrate a set of slices $\{\hat{\sigma}_{T_{i}}(K;\overline{p}_{i}^{*},N_{q}):i\leq N\}$, which are arbitrage-free in strike direction, and suppose that there is no calendar-spread arbitrage in the market. This means that, if $T_{i}\leq T_{j}$, then
@@
It is well-known [gatheral2014arbitrage] that the absence of calendar-spread arbitrage is equivalent to the condition that the total implied variance is increasing, i.e.
@@
Under this condition, a linear interpolation in total implied variance suffices to guarantee the absence of calendar arbitrage. Let $a\in[0,1]$ be such that $T=(1-a)T_{i}+aT_{j}$ and define
@@
Since the interpolation in time does not guarantee the absence of butterfly arbitrage in the interpolated slices, each interpolated slice must be checked separately for arbitrage in the strike direction.
@@ 4 Randomized Spot Volatility Parametrizations & Near Expiry Options for Equities
In this section of this paper, we consider an extension of the randomization of parametric volatility surfaces to a randomization of the “spot” parameter of the Black-Scholes formula, rather than implied volatility surface parameters. This new parametrization follows the same principles as the regular randomization described above, and is also defined as an expectation of a randomized price surface. The spot randomization replaces the input of the spot parameter in the Black-Scholes formula with a random variable $\vartheta$, which is centered at the current spot price $S_{0}$. We show below that, under certain conditions, such a randomization of the spot parameter is arbitrage-free, and that this parametrization is particularly effective to model option markets of very short-term options, known as near-expiry options. Near expiry options, also known as zero-day-expiry options (0DTE) options [^10] To be precise, 0DTE options have a time-to-maturity of less than a day. Since we also consider options of 1-3 days expiry, we prefer the expression near expiry options. are options whose maturity date is imminent, i.e., only a few days in the future, or even expire at the end of the day. On the day of an earning announcement, these options chains sometimes exhibit volatility slices with one or more concave sections, such as the W-shaped implied volatility slice or the “mustache”- shape [alexiou2023pricing, glasserman2023w]. The unusual shape of implied volatility is an indication of a bimodal or multimodal risk-neutral probability distribution [^11] Although note that this is neither a sufficient nor a necessary condition., i.e., a probability distribution function that exhibits two modes. The rationale is that the uncertainty from the earnings yields a multimodal risk-neutral probability distribution of $S_{T}$, reflecting the evolution of the stock price given different scenarios from the earnings (i.e., positive surprise vs negative surprise). Traditional implied volatility parametrizations, which have their origins in stochastic diffusion models, such as the SABR parametrization or the SVI, struggle to produce such shapes since the stochastic driver, which is a diffusion process, is inherently single-modal. The novel randomization method on the spot parameter offers a valuable solution in this case, as the randomization of the spot parameter offers an effective way to create superpositions of the PDFs with varying modes. This procedure yields a mixture density of a similar kind to well-established lognormal mixture model with varying modes [rebonato2004unconstrained, bloch2011smiling, glasserman2023w], which were developed as an extension of the classical lognormal mixture (as introduced by Brigo et al. [brigo2002lognormal]), where means of the mixed densities are not the same. We refer to these types of lognormal mixture parametrizations as multimodal lognormal mixture (MM LNM) parametrizations.
In this section we prove that the spot randomized volatility shape is arbitrage-free and show that it produces excellent results on the empirical data, utilizing data from the ticker AMZN on the day of the earnings announcement for Q1 in April 2018. The results show that the spot randomization has a better fit than the MM LNM parametrization, and a significantly better fit than the classical diffusion-based parametrizations SABR and SVI for this particular type of problem. Since the options we consider are short-expiry options, we focus the fitting of the data on a single volatility slice for a fixed time $T$. Although the randomization still provides an entire implied volatility surface, the main purpose of the spot randomization is to fit the non-standard volatility shape and to provide a viable price for any strike for the given expiry.
4.1 Randomized Spot Parameter